# Towards Enterprise-Ready Computer Using Generalist Agent

### v2

@@ Abstract
This paper presents our ongoing work toward developing an enterprise-ready Computer Using Generalist Agent (CUGA) system. Our research highlights the evolutionary nature of building agentic systems suitable for enterprise environments. By integrating state-of-the-art agentic AI techniques with a systematic approach to iterative evaluation, analysis, and refinement, we have achieved rapid and cost-effective performance gains, notably reaching a new state-of-the-art performance on the WebArena benchmark. We detail our development roadmap, the methodology and tools that facilitated rapid learning from failures and continuous system refinement, and discuss key lessons learned and future challenges for enterprise adoption.
@@ 1 Introduction
The development of enterprise-ready, computer-using generalist agents represents a significant frontier in artificial intelligence, poised to revolutionize productivity, workflows, automation, and decision-making across diverse industries. Recent advances in large language, vision, and action models, coupled with progress in agentic AI frameworks and implementations, are continuously raising the bar on existing computer-using benchmarks. While recent announcements, such as Anthropic’s Computer Use [1] and OpenAI’s Operator [4], suggest a growing commercial opportunity, realizing this vision requires more than just cutting-edge models, algorithms, or product prototypes, and significant challenges still remain.
At IBM Research, our ambition is to pioneer the development of agent systems that transcend mere task completion, and encompass the full spectrum of dimensions required for enterprise adoption, such as privacy, safety, trustworthiness, and cost-effectiveness of AI agentic solutions. As part of this mission, we have begun to develop a Computer Using Generalist Agent (CUGA). Our vision for IBM CUGA is to develop a generalist agent that can be adapted and configured by knowledge workers to perform routine or complex aspects of their work in a safe and trustworthy manner. Our first version focuses on knowledge worker tasks within web applications, and we tested it on the WebArena benchmark [7]. On WebArena, our agent achieves a new state of the art result of 61.7% on task completion [6].
This paper details the current state of our work, outlining the evolution of our agentic architecture to address the challenges posed by the WebArena benchmark. We describe our iterative development methodology and the tools that facilitated rapid learning from failures and cost-effective architectural improvements. Furthermore, we share lessons learned and highlight key challenges in realizing the full potential of such systems. Our primary contribution lies in disseminating the methodology, architecture, and practical experience that enabled us to achieve top performance on the WebArena leaderboard. Additionally, we have created a fully interactive dashboard  [^1] IBM CUGA dashboard - https://cuga.dev/ showcasing our performance results and agent trajectories, enhancing transparency.
@@ 2 Methodology and Tools
Our approach to developing an enterprise-ready computer using generalist agent (CUGA) is grounded in a philosophy of iterative evolution and rapid learning. We began with a simple agent architecture, intentionally designed to be a starting point, and committed to refining it based on empirical results and failure analysis. This evolutionary strategy allows us to adapt quickly to the complexities of real-world scenarios and continuously improve performance.
A cornerstone of our development methodology is the implementation of a smart sampling strategy. Recognizing the time-intensive nature of evaluating agent systems on comprehensive benchmarks, we adopted a technique of selecting an initially small, representative subset of the benchmark, enlarging the subsets as the system evolved to be better and more stable. This allowed us to rapidly test hypotheses, identify failure areas and side effects, and iterate on improvements before scaling up to larger portions of the benchmark. This approach significantly accelerates our learning cycle, enabling us to achieve rapid performance gains. Figure 1 depicts the main phases in our iterative evaluate-analyze-enhance process. In each iteration we evaluate on a larger sample, validate that expected performance gains are achieved, we analyze failures and prioritize areas of improvements that would maximize the performance gain in the following cycle.
Figure 1: The evaluate-analyze-enhance iterative development process.
@@
Figure 2: CUGA performance dashboard providing an overview and details performance results per task, with drill down links into trajectories
@@
    Comparative Analysis: Building upon the performance dashboard, this view enables direct comparison of results between different agent versions. It highlights previously resolved failures that are now successful, new failures on new data points, as well as failures on previously successful runs. This feature allows for rapid assessment of the impact of changes, validation of hypotheses, detection of side effects, and for regression purposes.
@@
    Trajectory Visualization and error classification: This tool, depicted in Figure 3 allows us to delve into individual failure cases, visualizing the agent’s interaction with the environment, its perception, reasoning process, and actions taken by different components. This detailed view enables us to quickly pinpoint and classify the root cause of failures and generate targeted hypotheses for improvement.
@@
Figure 3: The CUGA trajectory visualization, provides an easy access to the environment observation as screenshot, perception as accessibility tree, the action instruction, and its element grounding in the accessibility tree
@@ 3 Architecture Evolution
This section details the evolutionary journey of IBM CUGA’s agentic architecture. Our design philosophy centered on an iterative approach: beginning with a basic agent architecture and progressively enhancing it based on rigorous failure analysis and a prioritization of performance-maximizing improvements. Building upon this foundation, our system has evolved into a complex multi-agent architecture, leveraging LangGraph [^2] LangGraph framework - https://www.langchain.com/langgraph for managing stateful coordination between all the agents, and LangChain [^3] LangChain - https://www.langchain.com/ for common interface againts a mix of open and frontier LLM models. We use Playwright to control the browser, and the screenshot and accessibility tree for the observation space. We leverage the evaluation code from BrowserGym [2] to evaluate ourselves against WebArena. A simplified high-level representation of the final architecture is depicted in Figure 4. In the following subsections we further detail a chronicle of the key evolutionary cycles of the architecture.
Figure 4: A simplified high-level representation of IBM CUGA architecture, illustrating the interaction between user intent, context enrichment, high level plan controller, sub task plan-execute agents, environment action and observation, and learning and knowledge components.
@@ 3.1 Addressing Long-Horizon and Complex Tasks
Our initial iteration implemented a simple Plan-Act-Observe agentic loop and was evaluated on a small, representative sample of the WebArena dataset (3-5 sample templates per application domain). This baseline architecture achieved 15% task completion on this limited subset. However, it quickly became apparent that this approach was insufficient for handling complex, long-horizon tasks. These tasks often require the planner to orchestrate a sequence of actions and decisions while maintaining context regarding original goals and progress. Furthermore, the initial architecture struggled with tasks involving multi-site control or data flow, copy/paste operations, and the manipulation of lists and loops.
@@
    Plan Controller Agent: This agent is responsible for high-level planning, decomposing complex tasks into sub-tasks, selecting optimal sub task sequencing, handling loops and lists, tracking sub-task progress, and determining task completion.
@@
    Sub-task Plan-Execute Agents: These agents focus on the local planning of individual steps, UI element grounding (locating), and interaction.
This decomposition allowed for more robust handling of complex tasks, as the Plan Controller Agent could manage the overall task flow, while the Sub-task Plan-Execute Agents could handle the specifics of interacting with the UI. This separation of concerns significantly improved performance and laid the groundwork for further architectural refinements.
3.2 Enhanced Grounding, Interaction, and Observation
Building on the previous iteration, we refined our sampling strategy to create a larger and more representative sample of the WebArena benchmark. This expanded dataset, comprising 44 data points encompassing both previously encountered and newly identified task types, provided a more comprehensive evaluation platform. With this refined dataset, we achieved a task completion rate of 37.8%. However, failure analysis revealed persistent challenges for the Web application Sub-task Plan-Execute agents in the following areas:
@@ 3.2 Enhanced Grounding, Interaction, and Observation
    Action Execution Discrepancies: The planner often correctly identified the necessary action (e.g., selecting from a dropdown, typing into a search box), but the Action agent struggled to translate this into specific interaction steps. This issue frequently stemmed from the diverse and sometimes non-standard implementations of UI elements across different websites.
@@
    UI Element Grounding Failures: The Action agent, while attempting to perform an action, frequently failed to accurately locate (ground) the target UI element.
@@
    Complex Interaction and Extraction: Performing both UI element interaction and complex information extraction within a single agent’s prompt proved overly demanding. This necessitated a different perception mechanism and a separation of responsibilities.
@@
    Popup Obstruction: Pop-up windows occasionally obscured the agent’s observation space, hindering its ability to perceive the environment accurately.
To address these challenges, we implemented the following enhancements within the Browser Sub Task Agent:
@@
    Robust Element Interaction: We augmented the Action agent with an immediate feedback loop for element interaction. This feedback mechanism allows the agent to explore alternative interaction strategies and bypass irrelevant or obstructing UI elements (e.g. popups), improving its ability to handle diverse UI implementations.
@@
    Dedicated Information Extraction Agent: We introduced a specialized Information Extraction agent. This agent receives a distinct observation space, separate from the Action agent responsible for page interaction. This separation of concerns allows each agent to focus on its specific task, improving both interaction and extraction accuracy.
@@
    Enhanced Visual Context: We enriched the agent’s perception by incorporating screenshots in addition to the accessibility tree. This added visual context improves both the agent’s decision-making process for subsequent actions and its ability to ground UI elements.
@@ 3.3 Enhancing Stability and Mitigating Hallucinations
Large Language Models and agentic AI systems inherently exhibit variability in their execution. Running the same code on the same benchmark and environment can yield diverse reasoning paths and action sequences. While this characteristic fosters creative problem-solving, alternative exploration, and diverse decision-making, it can also lead to inconsistencies and hallucinations. To mitigate these issues and enhance stability, we incorporated reflection, critique, and judgment techniques within several of our agents. These techniques were implemented both through prompt engineering, refining existing prompts to encourage self-assessment, and by introducing dedicated reflection/judge agents. This dual approach allowed us to address both the underlying reasoning processes and the final outputs, improving the reliability and consistency of CUGA’s performance.
@@ 3.4 Planner Alignment through Context Enrichment, Learning, and Knowledge Injection
    Application Navigation Knowledge Acquisition: For each newly encountered application, this component explores the application’s navigation space, effectively mining a site-map-like structure. This structural knowledge enriches the context for intents that require a comprehensive understanding of the application’s functionality, guiding the planners toward appropriate actions.
@@
    Contextual Enrichment based on Task Assessment: Based on the assessed task, this component injects relevant application navigation knowledge and other contextual information, further refining the planners’ understanding of the task requirements and the available tools.

### v3

@@ Abstract
This paper presents our ongoing work toward developing an enterprise-ready Computer Using Generalist Agent (CUGA) system. Our research highlights the evolutionary nature of building agentic systems suitable for enterprise environments. By integrating state-of-the-art agentic AI techniques with a systematic approach to iterative evaluation, analysis, and refinement, we have achieved rapid and cost-effective performance gains, notably reaching a new state-of-the-art performance on the WebArena and AppWorld benchmarks. We detail our development roadmap, the methodology and tools that facilitated rapid learning from failures and continuous system refinement, and discuss key lessons learned and future challenges for enterprise adoption.
@@ 1 Introduction
The development of enterprise-ready, computer-using generalist agents represents a significant frontier in artificial intelligence, poised to revolutionize productivity, workflows, automation, and decision-making across diverse industries. Recent advances in large language, vision, and action models, coupled with progress in agentic AI frameworks and implementations, are continuously raising the bar on existing computer-using benchmarks. While recent announcements, such as Anthropic’s Computer Use [1] and OpenAI’s Operator [4], suggest a growing commercial opportunity, realizing this vision requires more than just cutting-edge models, algorithms, or product prototypes, and significant challenges remain.
At IBM Research, our ambition is to pioneer the development of agent systems that transcend mere task completion and encompass the full spectrum of dimensions required for enterprise adoption, such as privacy, safety, trustworthiness, and cost-effectiveness of AI agentic solutions. As part of this mission, we have begun to develop a Computer Using Generalist Agent (CUGA). Our vision for IBM CUGA is to develop a generalist agent that can be adapted and configured by knowledge workers to perform routine or complex aspects of their work in a safe and trustworthy manner. Our first version focuses on knowledge worker tasks within web applications, and we tested it on the WebArena benchmark [8].
Table 1: CUGA Results on Benchmarks
Benchmark Name Success/Completion Rate Domain Description
WebArena 61.7% success rate Web Evaluates web-based tasks
AppWorld 48.2% scenario completion rate API Evaluates API tasks
As shown in Table 1, CUGA achieves a new state-of-the-art result of 61.7% task completion on the WebArena benchmark [7]. On the AppWorld benchmark [5], which challenges agents to complete complex, multi-step workflows across diverse API-driven applications, the agent achieves a 46% scenario completion rate—also setting a new state-of-the-art. Notably, AppWorld evaluates an agent’s ability to dynamically select appropriate APIs, manage variables, reason about preconditions and outputs, and align its strategy with long-horizon goals—skills that are essential for real-world enterprise systems.
This paper outlines the current state of our work, detailing the evolution of our agentic architecture in response to the challenges posed by the WebArena and AppWorld benchmarks. We describe our iterative development methodology and the tools that facilitated rapid learning from failures and cost-effective architectural improvements. Furthermore, we share lessons learned and highlight key challenges in realizing the full potential of such systems. Our primary contribution lies in disseminating the methodology, architecture, and practical experience that enabled us to achieve top performance on both the WebArena and AppWorld leaderboards. Additionally, we have created a fully interactive dashboard [^1] IBM CUGA dashboard: https://cuga.dev/ that showcases our performance results and agent trajectories.
@@ 2 Methodology and Tools
Our approach to developing an enterprise-ready computer using generalist agent (CUGA) is grounded in a philosophy of iterative evolution and rapid learning. We began with a simple agent architecture, intentionally designed to be a starting point, and committed to refining it based on empirical results and failure analysis. This evolutionary strategy enables us to adapt quickly to the complexities of real-world scenarios and continually improve our performance.
A cornerstone of our development methodology is the implementation of a smart sampling strategy. Recognizing the time-intensive nature of evaluating agent systems on comprehensive benchmarks, we adopted a technique of selecting an initially small, representative subset of the benchmark and then enlarging the subset as the system evolved to become better and more stable. This allowed us to rapidly test hypotheses, identify failure areas and side effects, and iterate on improvements before scaling up to larger portions of the benchmark. This approach significantly accelerates our learning cycle, enabling us to achieve rapid performance gains. Figure 1 depicts the main phases in our iterative evaluate-analyze-enhance process. In each iteration, we evaluate the system on a larger sample, validate that expected performance gains are achieved, analyze failures, and prioritize areas of improvement that would maximize the performance gain in the following cycle.
Figure 1: CUGA’s iterative evaluate-analyze-enhance cycle, enabling rapid failure diagnosis, targeted improvements, and continuous performance gains.
@@
Figure 2: CUGA performance dashboard providing an overview and detailed performance results per task, with drill-down links into trajectories
@@
    Comparative Analysis: Building upon the performance dashboard, this view enables direct comparison of results between different agent versions. It highlights previously resolved failures that are now successful, new failures on new data points, as well as failures on previously successful runs. This feature enables the rapid assessment of the impact of changes, validation of hypotheses, detection of side effects, and facilitates regression analysis.
@@
    Trajectory Visualization and error classification: This tool, depicted in Figure LABEL:fig:cuga_trajectory allows us to delve into individual failure cases, visualizing the agent’s interaction with the environment, its perception, reasoning process, and actions taken by different components. This detailed view enables us to quickly pinpoint and classify the root cause of failures, generating targeted hypotheses for improvement.
@@
Figure 3: The CUGA trajectory visualization on the AppWorld task.
Figure 4: The CUGA trajectory visualization on the WebArena task provides easy access to the environment observation as a screenshot, perception as an accessibility tree, the action instruction, and its element grounding in the accessibility tree.
Figures 4 and 4: CUGA trajectory visualizations on API and web environments. Each illustrates the interaction between perception and action across modalities.
@@ 3 Architecture Evolution
This section details the evolutionary journey of IBM CUGA’s agentic architecture. Our design philosophy centered on an iterative approach: beginning with a basic agent architecture and progressively enhancing it through rigorous failure analysis and a focus on performance-maximizing improvements. Over time, this evolved into a sophisticated multi-agent system, composed of specialized sub-agents tailored for distinct interaction modalities: web and API. The web sub-agent is responsible for navigating and interacting with browser-based environments. It uses Playwright to control the browser and constructs its observation space from both screenshots and the accessibility tree. Evaluation against the WebArena benchmark is conducted using the official evaluation suite from BrowserGym [2]. The API sub-agent is designed to interact with structured application interfaces. It utilizes an API Registry to dynamically onboard and manage available applications by loading their OpenAPI schemas. This registry enables the agent to perform structured queries and hierarchical searches across APIs, scoped per application. Each application is backed by its own MCP server, which is automatically generated from its OpenAPI specification. This setup allows the agent to shortlist and invoke API actions in a scalable and semantically meaningful way. The entire system is orchestrated using LangGraph [^2] LangGraph framework - https://www.langchain.com/langgraph, which manages stateful coordination between agents, and LangChain [^3] LangChain - https://www.langchain.com/, which provides a unified interface for interacting with a mix of open and frontier LLMs. A simplified high-level representation of the final architecture is depicted in Figure 5. In the following subsections, we provide a detailed chronicle of the key evolutionary cycles of the architecture.
Figure 5: A simplified high-level representation of IBM CUGA architecture, illustrating the interaction between user intent, context enrichment, high-level plan controller, sub-task plan-execute agents, environment action and observation, and learning and knowledge components.
@@ 3.1 Addressing Long-Horizon and Complex Tasks
Our initial iteration implemented a simple Plan-Act-Observe agentic loop and was evaluated on a small, representative sample of the WebArena dataset (3–5 sample templates per application domain). This baseline architecture achieved a 15% task completion rate on this limited subset. We also evaluated the same architecture on a sample from the AppWorld benchmark training set, where it achieved only a 5% success rate. It quickly became apparent that this approach was insufficient for handling complex, long-horizon tasks. These tasks often require the planner to orchestrate a sequence of actions and decisions while maintaining context regarding original goals and progress. Furthermore, the initial architecture struggled with tasks involving multi-site control or data flow, copy/paste operations, and the manipulation of lists and loops. It also faced significant challenges in analyzing API responses, managing variables, and shortlisting relevant APIs for task execution.
@@
    Plan Controller Agent: This agent is responsible for high-level planning, decomposing complex tasks into sub-tasks, selecting optimal sub-task sequencing, handling loops and lists, tracking sub-task progress, determining task completion, and passing variables between sub-tasks to maintain continuity and context across the execution flow.
@@
    Sub-task Plan-Execute Agents: These agents focus on the local planning of individual steps, UI element grounding (locating), and interaction. There are two types of sub-task agents: browser-based agents, which interact with web interfaces, and API-based agents, which handle API calls, response parsing, and variable management.
This decomposition enabled more robust handling of complex tasks, as the Plan Controller Agent managed the overall task flow. At the same time, the Sub-task Plan-Execute Agents handled the specifics of interacting with either the user interface (UI) or APIs. This separation of concerns significantly improved performance and laid the groundwork for further architectural refinements.
3.2 Sub-Execution Agents Challenges and Enhancements
(a) This diagram outlines the architecture of an API Sub Agent, featuring an API Planner that coordinates task execution through short-term memory, strategic guidance, and variable management. It connects to two agents: the ShortlisterAgent, which selects applications and APIs via an MCP-Gateway, and the CodeAgent, which manages code tasks through a nested CodePlanner and CodeAgent that returns variables, enabling dynamic, context-aware decision-making across the system.
(b) This diagram presents the architecture of a Browser Sub Agent, structured around a central Browser Planner that includes a short-term memory (STM) for action history, a planning module, and a Reflection Judge for evaluating decisions. The system connects to a Select Agent and two specialized agents: the Action Agent, which executes browser interactions using tools such as Click, Type, Select, and Go Back; and the Question Answering Agent, which answers questions about the current page by processing web content through HTML-to-Markdown conversion and capturing screenshots. The architecture enables dynamic, memory-informed browsing and content extraction.
Figure 6: Architecture of the API and Browser Sub Agents
Building upon our prior iteration, we refined our sampling methodology to construct a more comprehensive and representative subset of the WebArena benchmark. This expanded dataset, comprising 44 data points spanning both previously encountered and novel task types, enabled a more rigorous evaluation. On this refined benchmark, our system achieved a task completion rate of 37.8%. However, detailed failure analysis revealed persistent limitations across both Browser and API Sub-task Plan-Execute agents.
3.2.1 Browser Sub-task Agent Challenges
@@
    Action Execution Discrepancies: While the planner frequently identified the correct high-level action (e.g., selecting from a dropdown, entering text), the Action agent often failed to translate this into precise interaction steps due to the heterogeneity and non-standard implementations of UI components across websites.
@@
    UI Element Grounding Failures: The agent frequently misidentified or failed to locate the intended UI element, resulting in execution errors.
@@
    Complex Interaction and Extraction: Combining UI interaction and complex information extraction within a single prompt proved overly burdensome, necessitating a decoupling of perception and action responsibilities.
@@
    Popup Obstruction: Pop-up elements occasionally occluded the observation space, impairing the agent’s ability to perceive and interact with the environment effectively.
3.2.2 Enhancements for Browser Sub-task Agent
@@
    Robust Element Interaction: We introduced a feedback loop within the Action agent to iteratively refine interaction strategies, enabling it to bypass obstructive or irrelevant UI elements (e.g., popups) and adapt to diverse UI implementations.
@@
    Dedicated Information Extraction Agent: A specialized agent was introduced to handle information extraction, operating on a distinct observation space. This separation improved both interaction fidelity and extraction accuracy.
@@
    Enhanced Visual Context: The agent’s perceptual input was augmented with screenshots alongside the accessibility tree, providing richer visual grounding for decision-making and UI element localization.
3.2.3 API Sub-task Agent Challenges
  1. 1.
    API Shortlisting Failures: The agent frequently failed to identify the most relevant API for a given sub-task, particularly in the presence of verbose or ambiguous OpenAPI specifications.
  2. 2.
    Variable Management Deficiencies: The agent lacked robust mechanisms for tracking and propagating variables across API calls and sub-tasks, leading to execution inconsistencies.
  3. 3.
    Unexpected Output Handling: The agent struggled to interpret and respond to unanticipated outputs from API calls or code execution, such as filtering or transforming intermediate results.
  4. 4.
    Lack of Global Planning Context: The agent operated in a myopic, step-wise fashion, lacking the ability to reason about global task structure or interdependent API workflows.
3.2.4 Enhancements for API Sub-task Agent
  1. 1.
    API Registry and Shortlisting: We developed an API registry that maps applications to their available APIs, enabling the agent to shortlist relevant APIs based on the sub-task and application context.
  2. 2.
    Concise OpenAPI Representation: To reduce token overhead and improve interpretability, OpenAPI specifications were transformed into a minimized, LLM-friendly format that preserved essential semantics while omitting extraneous metadata.
  3. 3.
    Variable Propagation Mechanism: A system-wide variable tracking framework was introduced, allowing the Plan Controller Agent to pass and update variables across sub-tasks, ensuring coherent multi-step execution.
  4. 4.
    Reflective Execution Loop: The API agent was augmented with a reflection mechanism, enabling it to re-evaluate and revise its strategy in response to unexpected outputs or execution failures.
@@ 3.3 Enhancing Stability and Mitigating Hallucinations
Large Language Models and agentic AI systems inherently exhibit variability in their execution. Running the same code on the same benchmark and environment can yield diverse reasoning paths and action sequences. While this characteristic fosters creative problem-solving, alternative exploration, and diverse decision-making, it can also lead to inconsistencies and hallucinations. To mitigate these issues and enhance stability, we incorporated reflection, critique, and judgment techniques within several of our agents. These techniques were implemented through both prompt engineering, which involved refining existing prompts to encourage self-assessment, and the introduction of dedicated reflection and judgment agents. This dual approach allowed us to address both the underlying reasoning processes and the final outputs, improving the reliability and consistency of CUGA’s performance.
@@ 3.4 Planner Alignment through Context Enrichment, Learning, and Knowledge Injection
    Application Navigation Knowledge Acquisition (for browser sub-agent): For each newly encountered application, this component explores the application’s navigation space, effectively mining a site-map-like structure. This structural knowledge enriches the context for intents that require a comprehensive understanding of the application’s functionality, guiding the planners toward appropriate actions.
@@
    Contextual Enrichment based on Task Assessment (for browser sub-agent): Based on the assessed task, this component injects relevant application navigation knowledge and other contextual information, further refining the planners’ understanding of the task requirements and the available tools.